OpenStudy (anonymous):

IF YOU HELP I WILL GIVE YOU A MEDAL AND FAN AND ANYTHING ELSE! The supplement of an angle is 24 degrees less than 4 times the complement of the angle. Find the measure of the angle.

OpenStudy (anonymous):

all you need to do is subtract

OpenStudy (anonymous):

lol ^^^true

OpenStudy (anonymous):

i need the equation

OpenStudy (anonymous):

24-4=you figure it out :D

OpenStudy (anonymous):

AHAHHAHAAHA

OpenStudy (anonymous):

can you guys please leave this is not helpful at all

OpenStudy (anonymous):

@Luigi0210 can you block them please

OpenStudy (anonymous):

me too i thought i was helping or are you just looking for the answer

OpenStudy (anonymous):

they told you what to do. you literally just need to subtract? why are you getting mad. wtf.

OpenStudy (anonymous):

saying you subtract really does not help. i need help finding the equation so either you stop making me feel stupid or you leave

OpenStudy (anonymous):

wait me too

OpenStudy (anonymous):

THE EQUATION IS TO SUBTRACT. lol girl. we ARE helping you.

OpenStudy (anonymous):

@stassiexoxo

OpenStudy (anonymous):

can you write out the equation

OpenStudy (anonymous):

24-4.

OpenStudy (blues):

What is the trouble here?

OpenStudy (anonymous):

there. lol. there ya go. can you figure it out?

OpenStudy (anonymous):

@blues i am asking for help and they just come on here without helping making me feel stupid

OpenStudy (anonymous):

@kendallnicole thats actually not the equation so leave

OpenStudy (anonymous):

oooooh girly over here's gettin feisty. byeee. don't say we didn't help ya.

OpenStudy (anonymous):

can someone please just write out the equation for me

OpenStudy (anonymous):

lol funny you didnt @kendallnicole

OpenStudy (anonymous):

@kendallnicole why are you still here? leave please

OpenStudy (anonymous):

so what is the question again

OpenStudy (texaschic101):

180 - x = 4(90 - x) - 24 180 - x = 360 - 4x - 24 180 - x = 336 - 4x 180 - 336 = -4x + x - 156 = -3x -156/-3 = x 52 = x check.. 180 - x = 4(90 - x) - 24 180 - 52 = 4(90 - 52) - 24 128 = 4(38) - 24 128 = 152 - 24 128 = 128 (correct) angle is 52 degrees

OpenStudy (anonymous):

^ she just helped

OpenStudy (anonymous):

@texaschic101 thank you so much!!!!!

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

sorry i made you feel that way

OpenStudy (anonymous):

can you please leave i have way too much on my mind and i dont want this on top of everything.

OpenStudy (luigi0210):

It's okay guys, just move it along.

OpenStudy (texaschic101):

she just apologized....that was the right thing to do

OpenStudy (anonymous):

i was saying sorry why are you being mean

OpenStudy (anonymous):

ok thank you for apologizing

OpenStudy (anonymous):

sorry for being mean

OpenStudy (anonymous):

ok now i can leave i was getting desperate

OpenStudy (anonymous):

i almost brought the puppy face LOL

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

@texaschic101 thank you again

OpenStudy (texaschic101):

glad to help :)